OpenStudy (anonymous):

2x+3y=6 3x-y=-13, solve by substitution or elimination method. can someone show me step by step how to solve this

OpenStudy (anonymous):

u have flvs

OpenStudy (anonymous):

do i have what?

OpenStudy (anonymous):

never mind is a online school

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

2x+3y=6 3x-y=-13 multiply the second equation by 3 you get 9x-3y=-13 you add the first equation to your new equation and you get 11x=-33 solve for x x=-3 substitute the new x to your first equation and you get -6 + 3y=6 3y=12 y=4 check your answer 2(-3)+3(4)=6 -6+12=6 6=6